OpenStudy (anonymous):

soybean meal is 18% protein; cornmeal is 9% protein . How many pounds of each should be mixed together in order to get 360lb mixture thats 10% protein

OpenStudy (anonymous):

We need to equations, the first one deals with pounds s + c = 360

OpenStudy (anonymous):

The second equation deals with protein soybean protein + cornmeal protein = mixture protein .18s + .09c = .10(s + c) or .18s + .09c = .10(360)

OpenStudy (anonymous):

Do you know how to solve s + c = 360 .18s + .09c = .10(360) after multiplying the second equation by 10 s + c = 360 18s + 9c = 3600 Would you use elimination or substitution?

OpenStudy (anonymous):

substitution I think maybe im wrong

OpenStudy (anonymous):

I got the 3600 in my math but i got lost after that

OpenStudy (anonymous):

\[.18x+.09(360-x)=36\]solve for x

OpenStudy (anonymous):

probably easier to get rid of annoying decimals and solve \[18x+9(360-x)=3600\]

OpenStudy (anonymous):

yes i got rid of the decimals also

OpenStudy (anonymous):

\[18x+3240-9x=3600\] \[9x=3600-3240\] \[9x=360\] \[x=360\div9=40\]

OpenStudy (anonymous):

so now would i do the same but put 18 in the place of 9